# Statistical significance in choice modelling: computation, usage and reporting

### v2

@@ Abstract
This paper offers a commentary on the use of notions of statistical significance in choice modelling. We argue that, as in many other areas of science, there is an over-reliance on 95% confidence levels, and misunderstandings of the meaning of significance. We also observe a lack of precision in the reporting of measures of uncertainty in many studies, especially when using $p$-values and even more so with star measures. The paper provides a precise discussion on the computation of measures of uncertainty and confidence intervals, discusses the use of statistical tests, and also stresses the importance of considering behavioural or policy significance in addition to statistical significance.
@@ 1 Introduction
There is growing criticism of the over-reliance on statistical significance, and its misuse, in several branches of science, calling for care in interpretation (see e.g. Wasserstein and Lazar, 2016), the use of much stricter tests (Benjamin et al., 2017), or even a complete abandonment of the notion of statistical significance (Amrhein et al., 2018; Wasserstein et al., 2019). Against this background, the timing seems apt to offer a commentary on this issue within the field of choice modelling.
We focus on four key topics, namely:
  • the computation of measures of uncertainty, i.e. standard errors;
  • their use in statistical tests and analyst decisions on model specification;
  • the computation and use of confidence intervals; and
  • the reporting of uncertainty and test outcomes.
@@
In our discussions, we question the strict adherence to the 95% level of confidence, highlight misconceptions as to what this implies, and highlight a lack of precision in some reporting practices. We also stress the need for more precise language in describing the outcome of significance tests and further call for analysts to not focus on statistical significance alone, but also consider the importance of a finding from a behavioural or policy perspective. While the points raised in this paper will be obvious to many choice modellers with a strong econometric background, this seems to not be the case in the more applied community. This conclusion is supported by the recent discussions in Parady and Axhausen (2023) who note a lack of rigour in the reporting of results and their statistical properties.
The remainder of this paper is organised as follows. Section 2 discusses the reasons for measures of uncertainty and their calculation. This is followed in Section 3 by a focus on confidence intervals, and Section 4 on statistical tests. Section 5 presents an empirical example, before Section 6 offers some conclusions.
@@ 2 Measures of uncertainty
Let us assume that we have a sample of $N$ people [^1] Here, $N$ relates to the sample actually used in estimation, i.e. excluding any hold out data set aside for validation.. Let us further denote the chosen alternative [^2] In this simple illustration, we focus on a cross-sectional context, with one choice situation per person. Extensions to repeated choice data are straightforward, with $Y_{n}$ simply being a sequence of choices. for person $n$ as $Y_{n}$. The analyst specifies a model that uses a vector of model parameters $\beta$ (of length $K$), with $\beta=\left<\beta_{1},\ldots,\beta_{K}\right>$. The log-likelihood function for this model is then given by:
@@
[…]
The frequentist (or classical) view holds that, for each single parameter $\beta_{k}$, there exists, subject to the chosen model specification, a unique true (or population) value $\beta_{k}^{*}$, but that our data is incomplete (it is a sample) and we thus have sampling error in our estimate $\hat{\beta}_{k}$ for parameter $\beta_{k}$. An estimate without uncertainty would thus only be obtained if the entire population was sampled. It should already be noted that this theory also relies on the conditions that 1) the model is correctly specified, and 2) that the model is identified, i.e. not overspecified, thus for example requiring appropriate normalisations.
In practice, an analyst would of course not repeat the sampling process to build the sampling distribution, but would estimate the model on the entire set of collected data, and would also obtain, from the same sample, estimators $\hat{\mu}_{\beta_{k}}$ and $\hat{\sigma}_{\beta_{k}}$ of the parameters of the sampling distribution.
For DCMs, we can only make asymptotic statements about the sampling distribution, that is, statements that relate to a case when the sample size $N$ goes to infinity. If a DCM is correctly specified, then, as sample size $N$ increases, the MLE for our vector $\beta$, i.e. $\hat{\beta}$, converges to a normal distribution [^3] We refer the reader to Cameron and Trivedi (2005, chapter 5). around the vector of true population values $\beta^{*}$, with:
@@
Under mild regularity conditions (see e.g. Lehmann and Casella, 1998, Lemma 5.3), it holds that $I=-A$, such that $\Omega=\frac{1}{N}A^{-1}IA^{-1}$ simplifies to $\Omega=\frac{1}{N}I^{-1}$ or $\Omega=-\frac{1}{N}A^{-1}$. To determine a sample estimate of the asymptotic variance matrix evaluated at $\hat{\beta}$, denoted by $\hat{\Omega}$, one can compute $\hat{A}$ and $\hat{I}$,
@@
where $\hat{H}$ is the Hessian and $\hat{O}$ is the sum of outer products of partial derivatives, both evaluated in the estimates $\hat{\beta}$. We can compute three different estimates of the variance matrix as shown in Table 1. The “classical” estimate of the variance matrix is based on the negative inverse of the Hessian, while the robust variance matrix is calculated by “sandwiching” $\hat{O}$ between $\hat{H}^{-1}$. Finally, the BHHH estimate (Berndt et al., 1974) may be attractive for computational reasons, but may differ substantially from the classical estimate.
@@
Berndt–Hall–Hall-Hausman (BHHH) $\hat{\Omega}=\textstyle{\frac{1}{N}}\,\hat{I}^{-1}\;=\;\hat{O}^{-1}$
Robust sandwich $\hat{\Omega}=\textstyle{\frac{1}{N}}\,\hat{A}^{-1}\,\hat{I}\,\hat{A}^{-1}\;=\;\hat{H}^{-1}\,\hat{O}\,\hat{H}^{-1}$
@@
If the model is correctly specified, and the optimum is well defined, the sampling variance of the MLE, i.e., $\hat{\Omega}$, will correspond to the Cramér-Rao lower bound, i.e. the estimates are minimum variance. From $\hat{\Omega}$, we can obtain estimated standard errors as the square roots of its diagonal elements, say $\hat{\sigma}_{k}=\sqrt{\hat{\Omega}[k,k]}$ for parameter $\beta_{k}$. As discussed in detail by Daly et al. (2012), error measures can similarly be computed for functions of different model parameters, such as differences or ratios, using the Delta method, with the same underlying MLE properties. Our discussions in terms of confidence intervals and hypothesis testing thus apply in the same way to functions of parameters (e.g., willingness-to-pay) as to individual estimates.
@@
A number of subtleties arise here. First, these properties rely on the notion that the model itself is correctly specified, which is in most cases a dubious claim to make at best. For example, if an analyst uses a linear specification when the true specification is non-linear, then the true values for our model parameters could not be obtained due to endogeneity (Guevara, 2024). By using ML estimation, the analyst will obtain consistent estimates of the parameters for the specified model, but these estimates, and their standard errors, will in general be affected by endogeneity if the specified model is not the correct model. In other words, the wrongly specified model has wrong estimates and wrong statistical significance properties. The standard error for $\hat{\beta}$ simply reflects the uncertainty in our estimated values for the estimated model, and will in general be wrongly estimated also. They are not related to any inconsistency caused by choosing the wrong model. This already means that a wrongly specified model can lead to wrong inferences about statistical significance. One specific example of misspecification is the inclusion of irrelevant, or exclusion of relevant (i.e. missing), variables. This kind of misspecification not only influences the parameter estimates and potentially causes inconsistency, but also affects the size of the standard errors of all the parameters in the model due to its impact on the information matrix. Second, the properties relate to a distribution of the MLE around the true parameter value for the population, not for the sample. The assumption that the model estimation yields the true value for the parameter for the sample, conditional on the model specification, is implicit, i.e. disregarding the possibility of convergence to an inferior local optimum that would give a inconsistent estimate of $\beta$ for the current model and sample.
@@
We already mentioned that the above calculations rely on the condition that the model is correctly specified. However, under the mild specification mistakes that do not compromise consistency, such as in some cases neglected heterogeneity or correlation between observations, two key approaches exist that make fewer assumptions than the classical covariance matrix calculations and thus provide a less conservative estimate of the sampling error.
@@
The first approach is to use robust standard errors, which, as discussed above, are obtained using the so called sandwich estimator [^4] The origin of the term sandwich estimator is that $\Omega$ forms the bread, while the BHHH matrix is the filling.. Daly et al. (2023) show that the Delta method can also be applied using the sandwich matrix to obtain robust estimates of errors in functions of estimated parameters.
@@
If the model is correctly specified, asymptotically $\hat{\Omega}_{B}=\hat{\Omega}$, in which case the robust and classical covariance matrices are equal. One typical implementation is that the summation in the BHHH matrix is across individual people rather than choice observations, i.e. as in Equation (7). This ensures that the calculation accounts for correlation across choices for the same person that the modeller may have omitted from the model specification, recognising that $T$ observations each from $N$ people is not the same as $1$ observation each from $NT$ people. In such a case with repeated choice data, one can expect robust standard errors to be larger than classical standard errors because of the inherent correlation in choices across individuals not recognised by the classical standard errors. Our personal experience, however, is that robust standard errors can be both larger and smaller than classical standard errors, and that larger differences for specific parameters indicate a higher degree of misspecification for these parameters. Specifically, when robust standard errors are smaller than classical standard errors, this tends to indicate substantive misspecification or identification issues. Unfortunately, we do not have analytical proofs for these observations, but we refer the interested reader to King and Roberts (2015) to learn more about the extent to which robust standard errors identify methodological problems (but do not fix them).
@@
The second approach is to use non-parametric methods, like bootstrapping. This works by repeated sampling, and is the approach with least assumptions, but is computationally most demanding. Bootstrapping recognises that the source of uncertainty is sampling error, and approximates this by simulating the sampling process with the data at hand. In particular, we would draw $S$ versions of the data with replacement (i.e. some people will be included multiple times) from the original sample, producing samples of the same size as the original data. The specified model would then be estimated on each of the samples, yielding $S$ sets of estimates of $\beta$, say $\hat{\beta}^{\left(s\right)}=\left<\hat{\beta}_{1}^{\left(s\right)},\ldots,\hat{\beta}_{K}^{\left(s\right)}\right>,\,s=1,\ldots,S$. If $S$ is sufficiently large, we obtain the result that:
@@
By performing the bootstrapping at the level of individual decision-makers as opposed to individual choice observations, our bootstrap standard error also recognises the repeated choice nature of the data, in line with the robust standard errors using the BHHH matrix at the person level in Equation (7).

### v3

@@ Abstract
This paper offers a commentary on the use of notions of statistical significance in choice modelling. We review the reasons for uncertainty in parameter estimates, provide a precise discussion on the computation of measures of uncertainty and confidence intervals, and discuss the use of statistical tests. We argue that, as in many other areas of science, there is an over-reliance on 95% confidence levels, and misunderstandings of the meaning of significance. We also observe a lack of precision in the reporting of measures of uncertainty in many studies, especially when using $p$-values and even more so with star measures. The paper also stresses the importance of considering behavioural or policy significance in addition to statistical significance. Finally, we stress a number of points that are specific to choice modelling and which require special attention, notably in relation to derived measures such as willingness-to-pay, the treatment of random heterogeneity, and the use of repeated choice data.
@@ 1 Introduction
Empirical work focusses on producing metrics of interest, with a particular emphasis on investigating the existence of specific effects or phenomena. In this context, analysts often formulate hypotheses, and then run tests to establish whether an effect exists. When using mathematical structures (models) to represent the data, tests also look at whether one model explains the data better than another. With finite and limited data, any empirical findings have uncertainty attached to them, and the recognition of this leads analysts to attach measures of confidence to their findings, under the broad umbrella of “statistical significance”.
@@
While ubiquitous across different methodological disciplines and fields of application, there is also growing criticism of the over-reliance on statistical significance, and its misuse, in several branches of science. This is exemplified in the wide ranging discussions by Ziliak and McCloskey (2008), or more recent papers calling for care in interpretation (see e.g. Wasserstein and Lazar, 2016), the use of much stricter tests (Benjamin et al., 2017), or even a complete abandonment of the notion of statistical significance (Amrhein et al., 2018; Wasserstein et al., 2019).
A key component of this criticisms relates to excessive attention to the question whether an effect exists (i.e. is it different from zero) rather than whether an effect matters (i.e. how large it is). Similarly, there is misunderstanding about what statistical tests tell us, as highlighted in the discussions by Ziliak and McCloskey (2008) on the “fallacy of the transposed conditional”, with a clear distinction between the probability of the evidence given a hypothesis and the probability of the hypothesis given the evidence.
Against this background, the timing seems apt to offer a commentary on this issue within the field of choice modelling. A number of specific considerations arise in our field, making a separate discussion appropriate. We list six of these reasons in what is a non-exhaustive list:
  1. 1.
    Choice modelling relies on the results of a model for which parameters are estimated, meaning that measures of uncertainty cannot simply be calculated on the basis of the observed data. Instead, they rely on statistical theory, in particular in relation to the properties of maximum likelihood estimates (MLE), or the assumed data generating process and selected priors in Bayesian analysis. That is, also in Bayesian analysis, uncertainty is present, and even more specific.
  2. 2.
    Estimates from choice models have a different interpretation from for example estimates of a linear regression model, and this means that analysts rely on measures calculated on the basis of parameter transformations for actual insights, such as marginal rates of substitution. Other insights are obtained using model application, such as forecasting, especially if an analyst wants to establish the size of an effect beyond its existence per se. These outputs again need measures of uncertainty.
  3. 3.
    Applied choice modelling relies on the recognition and flexible treatment of extensive levels of heterogeneity in underlying preferences, leading to a requirement for a distinction between parameter uncertainty and distribution of preferences in a sample population.
  4. 4.
    There is a strong reliance on panel (or repeated choice) data, requiring either an explicit recognition of correlation across choices for the same person, or ex-post correction approaches to measures of uncertainty computed based on statistical theory developed for cross-sectional data.
  5. 5.
    A key focus in choice modelling relates to tests of structural assumptions of the underlying models, with extensive model comparison work where a number of tests have been developed for choice models in particular.
  6. 6.
    Finally, in addition to these field-specific considerations, it should also be recognised that the rapid expansion of choice modelling over the last couple of decades has opened the method up to analysts from a diverse set of backgrounds, including those with limited econometric training. The key underlying statistical concepts remain poorly understood by some, and statistical test and measures of uncertainty are used and reported inappropriately by many.
An important question arises before proceeding, namely why write a paper on significance measures and testing in the face of growing concerns about the usefulness of such tests? For example, Imbens (2021) argues that “…researchers are interested in a point estimate and the degree of uncertainty associated with that point estimate as the precursor to making a decision or recommendation to implement a new policy. In such cases, the absence or presence of statistical significance (in the sense of being able to reject the null hypothesis of zero effect at conventional levels) is not relevant”. Even more strongly, there are strong arguments against the usefulness of a hypothesis of a zero effect. Abadie (2020) state that “in economics … there are rarely reasons to put substantial prior probability on a point null” while Gelman et al. (2020) state that “… we do not generally think null hypotheses can be true: in social science and public health, just about every treatment one might consider will have some effect, and no comparison or regression coefficient of interest will be exactly zero.”
First, the usefulness of measures of uncertainty extends beyond statistical tests to the computation of confidence intervals, and for this, the background on computation is crucial. Second, we are of the opinion that notwithstanding the validity of some of these concerns, choice modellers will continue relying on statistical tests in this way, and our note at least serves to improve the way in which they do this, for example using one-sided tests in many cases.
@@
Our paper addresses the topic along a number of different dimensions. To provide a stronger foundation for understanding the notion of uncertainty, Section 2 discusses the reasons for measures of uncertainty and their calculation. The meaning of uncertainty is then explained further in the context of confidence intervals in Section 3. The central part of the paper, namely Section 4, focusses on statistical tests. We discuss the formation and testing of hypothesis, separately for individual results and for model selection, before turning to the reporting of uncertainty and test outcomes. In our discussions, we question the strict adherence to the 95% level of confidence, highlight misconceptions as to what this implies, and highlight a lack of precision in some reporting practices. We also stress the need for more precise language in describing the outcome of significance tests. In this context, we also caution analysts to pay attention to the size of effects rather than the existence of effects alone, thus also considering the importance of a finding from a behavioural or policy perspective. While the points raised in this paper will be obvious to many choice modellers with a strong econometric background, this seems to not be the case in the more applied community. This conclusion is supported by the recent discussions in Parady and Axhausen (2024) who note a lack of rigour in the reporting of results and their statistical properties. To support the theoretical discussions, Section 5 presents an empirical example, before Section 6 offers some conclusions.
@@ 2 Measures of uncertainty
2.1 Sampling error and parameter uncertainty
Our specific interest relates to model parameters (and later transformations thereof). The frequentist (or classical) view holds that, for each single parameter $\beta_{k}$, there exists, subject to the chosen model specification, a unique true (or population) value $\beta_{k}^{*}$, but that our data is incomplete (it is a sample) and we thus have sampling error in our estimate $\hat{\beta}_{k}$ for parameter $\beta_{k}$. An estimate without uncertainty would thus only be obtained if the entire population was sampled.
In practice, an analyst would of course not repeat the sampling process to build the sampling distribution, but would estimate the model on the entire set of collected data, and would also obtain, from the same sample, estimators $\hat{\mu}_{\beta_{k}}$ and $\hat{\sigma}_{\beta_{k}}$ of the parameters of the sampling distribution. We next turn to how this is derived in the context of maximum likelihood estimation.
2.2 From maximum likelihood estimation to the asymptotic covariance matrix
@@
Let us assume that we have a sample of $N$ people [^1] Here, $N$ relates to the sample actually used in estimation, i.e. excluding any hold out data set aside for validation.. Let us further denote the chosen alternative for person $n$ as $Y_{n}$. The analyst specifies a model that uses a vector of model parameters $\beta$ (of length $K$), with $\beta=\left<\beta_{1},\ldots,\beta_{K}\right$>. The log-likelihood function for this model is then given by:
@@
$\hat{\beta}=\underset{\beta}{\arg\max}\,LL_{N}\left(\beta\right).$ (2)
For DCMs, we can only make asymptotic statements about the sampling distribution, that is, statements that relate to a case when the sample size $N$ goes to infinity. In addition, it should already be noted that this theory also relies on the conditions that 1) the model is correctly specified, and 2) that the model is identified, i.e. not overspecified, thus for example requiring appropriate normalisations.
Under these conditions, as sample size $N$ increases, the MLE for our vector $\beta$, i.e. $\hat{\beta}$, converges to a normal distribution [^2] We refer the reader to Cameron and Trivedi (2005, chapter 5). around the vector of true population values $\beta^{*}$, with:
@@
Under mild regularity conditions (see e.g. Lehmann and Casella, 1998, Lemma 5.3), it holds that I=AI=-A, such that $\Omega=\frac{1}{N}A^{-1}IA^{-1}$ simplifies to $\Omega=\frac{1}{N}I^{-1}$ or $\Omega=-\frac{1}{N}A^{-1}$. To determine a sample estimate of the asymptotic variance matrix evaluated at $\hat{\beta}$, denoted by $\hat{\Omega}$, one can compute $\hat{A}$ and $\hat{I}$,
@@
where $\hat{H}$ is the Hessian and $\hat{O}$ is the sum of outer products of partial derivatives, both evaluated in the estimates $\hat{\beta}$. We can compute three different estimates of the variance matrix as shown in Table 1. The “classical” estimate of the variance matrix is based on the negative inverse of the Hessian, while the robust variance matrix is calculated by “sandwiching” $\hat{O}$ between $\hat{H}^{-1}$. Finally, the BHHH estimate (Berndt et al., 1974) may be attractive for computational reasons, but may differ substantially from the classical covariance matrix.
@@
Berndt–Hall–Hall-Hausman (BHHH) $\hat{\Omega}_{B}=\textstyle{\frac{1}{N}}\,\hat{I}^{-1}\;=\;\hat{O}^{-1}$
Robust sandwich $\hat{\Omega}_{R}=\textstyle{\frac{1}{N}}\,\hat{A}^{-1}\,\hat{I}\,\hat{A}^{-1}\;=\;\hat{H}^{-1}\,\hat{O}\,\hat{H}^{-1}=\hat{\Omega}\;\hat{\Omega}_{B}^{-1}\;\hat{\Omega}$
@@
If the model is correctly specified, and the optimum is well defined, the sampling variance of the MLE, i.e., $\hat{\Omega}$, will correspond to the Cramér-Rao lower bound, i.e. the estimates are minimum variance. From $\hat{\Omega}$, we can obtain estimated standard errors as the square roots of its diagonal elements, say $\hat{\sigma}_{k}=\sqrt{\hat{\Omega}[k,k]}$ for parameter $\beta_{k}$.
2.3 Robust covariance matrix and bootstrapping
@@
A number of subtleties arise in relation to the earlier discussions. First, the properties discussed above rely on the notion that the model itself is correctly specified, which is in most cases a dubious claim to make at best. For example, if an analyst uses a linear specification when the true specification is non-linear, then the true values for our model parameters could not be obtained due to endogeneity (Guevara, 2024). By using ML estimation, the analyst will obtain consistent estimates of the parameters for the specified model, but these estimates, and their standard errors, will in general be affected by endogeneity if the specified model is not the correct model. In other words, the wrongly specified model has wrong estimates and wrong statistical significance properties. The standard error for $\hat{\beta}$ simply reflects the uncertainty in our estimated values for the estimated model, and will in general be wrongly estimated also. They are not related to any inconsistency caused by choosing the wrong model. This already means that a wrongly specified model can lead to wrong inferences about statistical significance. One specific example of misspecification is the inclusion of irrelevant, or exclusion of relevant (i.e. missing), variables. This kind of misspecification not only influences the parameter estimates and potentially causes inconsistency, but also affects the size of the standard errors of all the parameters in the model due to its impact on the information matrix. Second, the properties relate to a distribution of the MLE around the true parameter value for the population, not for the sample. The assumption that the model estimation yields the true value for the parameter for the sample, conditional on the model specification, is implicit.
@@
Under mild specification mistakes that do not compromise consistency, such as in some cases neglected heterogeneity or correlation between observations, two key approaches exist that make fewer assumptions than the classical covariance matrix calculations and thus provide a less conservative estimate of the sampling error.
@@
The first approach is to use robust standard errors, which, as discussed above, are obtained using the so called sandwich estimator [^3] The origin of the term sandwich estimator is that $\Omega$ forms the bread, while the BHHH matrix is the filling., with:
$\hat{\Omega}_{R}=\hat{\Omega}\;\hat{\Omega}_{B}^{-1}\;\hat{\Omega}$, (8)
@@
where we refer the reader back to Table 2 for a definition of the individual components. If the model is correctly specified, asymptotically $\hat{\Omega}_{B}=\hat{\Omega}$, in which case the robust and classical covariance matrices are equal. In general, larger differences between classical and robust standard errors for specific parameters indicate a higher degree of misspecification for these parameters. Our personal experience is that, notwithstanding the general expectation that robust standard errors are larger than classical standard errors, the opposite can happen in some cases. Specifically, when robust standard errors are smaller than classical standard errors, this may indicate substantive misspecification or identification issues. Unfortunately, we do not have analytical proofs for these observations, but we refer the interested reader to King and Roberts (2015) to learn more about the extent to which robust standard errors identify methodological problems. Of course, it should be noted that robust standard errors do not fix these specification issues themselves, and this remains the task of the modeller. As the true model form is never known, it would thus be wise to continue using robust standard errors even after making improvements to the model on the basis of the discrepancy between classical and robust standard errors.
@@
The second approach is to use non-parametric methods, like bootstrapping. This works by repeated sampling, and is the approach with least assumptions, but is computationally most demanding, meaning it is rarely used, especially for complex models. Bootstrapping recognises that the source of uncertainty is sampling error, and approximates this by simulating the sampling process with the data at hand. In particular, we would draw $S$ versions of the data with replacement (i.e. some people will be included multiple times) from the original sample, producing samples of the same size as the original data. The specified model would then be estimated on each of the samples, yielding $S$ sets of estimates of $\beta$, say $\hat{\beta}^{\left(s\right)}=\left<\hat{\beta}_{1}^{\left(s\right)},\ldots,\hat{\beta}_{K}^{\left(s\right)}\right>,\,s=1,\ldots,S$. If $S$ is sufficiently large, we obtain the result that:
@@
2.4 Error measures for parameter transformations
While parameters from choice models provide initial insights into the studied behaviour, the key interest lies in transformations of such parameters. This includes most prominently marginal rates of substitution (MRS), such as willingness to pay (WTP). For simple linear specifications, these are given by ratios of parameters, while the computation becomes more difficult in the presence of non-linearity. Another reason for the need for transformation of parameters relates to the issue of effect sizes. Parameter values themselves do not tell us about the size of an effect in terms of impact on behavior, and neither do MRS measures, or indeed relative attribute importance (RAI). Instead, analysts need to look at demand predictions or elasticities (cf. Daly, 2024), or marginal effects (see e.g. Hess et al., 2022). Independently of which specific measure an analyst is thus using for further interpretation, it should be clear that these measures themselves are affected by uncertainty, given that they are derived on the basis of parameters with uncertainty.
As discussed in detail by Daly et al. (2012), error measures can be computed for any functions of different model parameters using the Delta method, with the same underlying MLE properties applying to those error measures. Daly et al. (2023) show that the Delta method can also be applied using the sandwich matrix to obtain robust estimates of errors in functions of estimated parameters. Our discussions in terms of confidence intervals in Section 3 and hypothesis testing in Section 4 thus apply in the same way to functions of parameters as to the estimates of the parameters themselves.
Additional complexities arise in the case of functions of randomly distributed coefficients. Let us assume that $\beta_{1}$ is allowed to vary randomly across individuals in a model. In the case of a continuous mixture model such as Mixed Logit, an analyst would obtain estimates of parameters that characterise the shape of the distribution of $\beta_{1}$, such as a mean and standard deviation. These parameters have uncertainty attached to them, thus also implying uncertainty in the extent of heterogeneity and shape of the distribution. We can then report uncertainty in these moments of the distribution, but this is not to confused with heterogeneity in $\beta_{1}$.
If we now turn to metrics computed on the basis of distributed coefficients, such as for example a MRS with $\frac{\beta_{1}}{\beta_{2}}$, when the numerator and denominator are distributed, the analyst needs to be particularly mindful of the difference between uncertainty, which relates to estimated parameters, and heterogeneity, which relates to the variation of sensitivities in a population. Model estimation provides parameters that characterise the shape of the latter, and these parameters have uncertainty attached to them, thus also implying uncertainty in the extent of heterogeneity and shape of the distribution.
In the case of both continuous mixture models such as Mixed Logit and discrete mixtures such as Latent Class, an analyst can compute moments such as mean and variance for the distribution of measures such as $\frac{\beta_{1}}{\beta_{2}}$. If these moments have a closed form expression, as they do in Latent Class and for some choices of distribution of $\beta_{1}$ and $\beta_{2}$ (such as Lognormal) for continuous mixtures, then the Delta Method can be applied to calculate measures of uncertainty for such moments. However, for distributions where the ratio $\frac{\beta_{1}}{\beta_{2}}$ does not have defined moments, or where an analyst wants to produce a measure of uncertainty around the entire continuous distribution of $\frac{\beta_{1}}{\beta_{2}}$, rather than just the moments thereof, more complex approaches are needed that involve a level of numerical simulation, as discussed in Bliemer and Rose (2013); Scaccia et al. (2023).
2.5 Repeated choice data
Our discussion so far has focussed on a cross-sectional context, with one choice situation per person. A large number, and possibly a majority, of applications of choice models however make use of repeated choice data, i.e. with multiple observations per individual, with likely correlation between them.
While a large share of studies explicitly model such correlations, for example by integrating over random distributions at the person (rather than observation) level (cf. Revelt and Train, 1998), this is not the case in work that relies on closed form models such as multinomial or nested logit. In such applications, post-estimation correction approaches are required. Indeed, the standard errors are likely to be underestimated, given that $T$ observations each from $N$ people is not the same as $1$ observation each from $NT$ people
Early correction approaches involved the use of Jackknife or Bootstrapping (cf Cirillo et al., 2000), but a more straightforward approach is to rely on a panel implementation of the sandwich or bootstrap covariance matrix.
For the sandwich approach, the typical implementation is for the summation in the BHHH matrix to be across individual people rather than choice observations, i.e. as in Equation (7). This ensures that the calculation accounts for correlation across choices for the same person that the modeller may have omitted from the model specification. In such a case with repeated choice data, one can expect robust standard errors to be larger than classical standard errors because of the inherent correlation in choices across individuals not recognised by the classical standard errors.
A similar correction is possible by performing bootstrapping at the level of individual decision-makers as opposed to individual choice observations. Analysts should be mindful that the use of either robust standard errors or bootstrapping does not imply that their model has captured the repeated choice nature of the data, but simply that a correction of the standard errors has been applied. It is of course always preferable to explicitly account for such correlations during model estimation, if empirically and computationally feasible.
2.6 Identification and over-specification issues